In each of the cases that follow, the components of a vector A⃗  are given. Use trigonometry to find the magnitude of that vector and the counterclockwise angle it makes with the +xaxis.

Ax = 3.00 m , Ay = 8.00 m

Please give your answer in the interval [0,360]∘.
Answer #1

magnitude,

A^2 = Ax^2 + Ay^2

A^2 = 3^2 + 8^2

A = 8.544 m

direction

phi = arctan ( 8/3)

phi = 69.44

=========

do comment in case any doubt, will reply for sure. Goodluck